what is the distance between the points (1,1) nad (-0.5,-1)

Found 2 solutions by stanbon, fcabanski:
Answer by stanbon(75887)

what is the distance between the points (1,1) and (-0.5,-1)
---
distance = sqrt[(1--1)^2 + (1--0.5)^2]
---
= sqrt[2^2+1.5^2]
----
= sqrt[4+2.25]
----
= sqrt[6.25]
----
= 2.5
==============
Cheers,
Stan H.
==============

Answer by fcabanski(1391)

Check my profile and contact me for one on one online tutoring.


The distance formula for the distance between two points is d = sqrt ((x2-x1)^2 + (y2-y1)^2).


In this case we'll call (1,1) point 1 and (-0.5,-1) point 2.


d = sqrt ((-.5-1)^2 + (-1-1)^2) = sqrt (2.25+4) = sqrt (6.25) = 2.5.